# Accelerated Mirror Descent for Non-Euclidean Star-convex Functions

### v1

@@ 4 Binary search for Hölder smooth functions
In our algorithm (Algorithm 1) we perform binary search to compute the momentum parameter $\lambda_{t}$ that satisfies
$\lambda g^{\prime}(\lambda_{t})+C_{t}g(\lambda_{t})\leq\epsilon_{t},$ (6)
where we use the abbreviation $g(\lambda)\triangleq F(\lambda x_{t}^{ag}+(1-\lambda)x_{t})-F(x_{t}^{ag})$. Condition (6) is later used in the convergence proof of Theorem 3.1. The core idea behind our Algorithm 2 is inspired by Algorithm 2 in [18]. The goal is to find a $\lambda_{t}$ that satisfies (6).
If $g(0)\leq g(1)$ or $g$ is decreasing at $\lambda=1$, then (6) is immediately satisfied. If neither of these conditions hold, then $g(0)>g(1)$ and $g^{\prime}(1)>0$, indicating that $g$ must switch from decreasing to increasing at some $\lambda_{*}\in(0,1)$, thus fulfilling $g^{\prime}(\lambda_{*})=0$.
We further show in the proof of Theorem 4.1 that all $\lambda\in[\lambda_{*}-\delta,\lambda_{*}]$ with a properly chosen $\delta$ also satisfy condition (6). The efficiency of our Algorithm in finding such a $\lambda$ is stated in the Theorem below.
Algorithm 2 Generalised binary search
@@
     if $g^{\prime}(\lambda_{t})+C_{t}g(\lambda_{t})\leq\epsilon_{t}$ then
@@
Under Assumption 1, particularly where $F$ is $(L,\kappa)$-weakly smooth, and for fixed $(C_{t},\epsilon_{t},x_{t},x_{t}^{ag})$ , the binary search algorithm finds $\lambda$, satisfying
@@
In [18], the binary search algorithm is analyzed for smooth functions, i.e., specifically with $\kappa=2$ in the weakly smooth sense. Our algorithm accommodates a general $\kappa$, addressing the broader class of weakly smooth functions. Although the running time depends on $\kappa$, it appears only as an exponent in the logarithm. Consequently, this results in a change by a constant factor, thereby generalizing the algorithm in [18].
Proof Sketch..
@@
The proof can be roughly divided into three parts. We first show that there exists a $\lambda_{*}$ satisfying the stronger condition (6) with $\epsilon_{t}=0$. Next, we show that this implies that condition (6) holds for all $\lambda\in[\lambda_{*}-\delta,\lambda_{*}]$, where $\delta$ is properly chosen depending on the weak smoothness parameters $(L,\kappa)$ of $F$ and the parameters $(C_{t},\epsilon_{t},x_{t},x_{t}^{ag})$ of the algorithm. Finally we prove that our algorithm finds a $\lambda$ within this interval within at most $\mathcal{O}\Big{(}\log\Big{(}\frac{LC_{t}\|x_{t}^{ag}-x_{t}^{md}\|^{\kappa}}{\epsilon_{t}}\Big{)}\Big{)}$ iterations. ∎
@@ 5 Conclusion and future research
In this work, we introduce a novel class of structured nonconvex functions, namely weakly smooth $\tau$-star-convex functions with respect to a norm satisfying the new regularity assumption introduced in Definition 2.7. We demonstrate that the regularity assumption is satisfied not only for any $p$-norm with $p>1$, but also for compositions of regular norms. Consequently, our results apply to a large subclass of star-convex functions, generalizing the setting of smooth star-convex functions with respect to the Euclidean norm [18].
We develop an accelerated algorithm (Algorithm 1) that efficiently minimizes any weakly smooth function within this broad class, extending the binary search technique from [18]. Our analysis shows that the algorithm achieves a near-optimal accelerated convergence rate for all $p$-norms with $p>1$. Additionally, our algorithm has a complexity dependence on $\tau^{\max\{2,p\}}$ for $p\geq 1$, which has been shown to be optimal in the Euclidean case when $p=2$ [18].
While this work focuses on unconstrained minimization problems, future research could extend our approach to constrained settings. Currently, the proximal step for calculating $x_{t+1}^{ag}$ in Algorithm 1 is valid only for unconstrained problems. The main challenge in constrained settings is evaluating the direction of the constrained proximal step.
Our convergence results rely on the assumption of uniform convexity of the distance-generating function $\psi$. While our analysis provides first insights into norms fulfilling this condition, such as (compositions of) $p$-norms, analysing different other regular norms that meet this criterion provides an opportunity for future research.
Finally, our method operates within a first-order access model, meaning that the algorithm has access to both the objective function value and gradient at the evaluation points. It remains an open problem whether this can be extended to “pure” first-order access models, that rely solely on the gradient evaluations. This can be done for the convex case, but it is not known whether this is possible for star-convex functions. Additionally, exploring the applicability of our method to a stochastic model presents another promising direction. Although stochastic accelerated mirror descent has been previously analyzed in the literature [19], the main challenge lies in adapting the binary search algorithm to the noise.
Appendix A Proofs for preliminaries
Recall the notion of $C$-regularity from Definition 2.7.
@@ Appendix A Proofs for preliminaries
For $p>1$, $\|\cdot\|_{p}$ is $1$-regular.
Proof.
We denote by $p_{\ast}$ the Hölder conjuguate of $p$, defined such that $\frac{1}{p}+\frac{1}{p_{\ast}}=1$. Recall that for $x\in\mathbb{R}^{d}$, $\|x\|_{p}:=(\sum_{i}|x_{i}|^{p})^{\frac{1}{p}}$, which corresponds to the Euclidean norm for $p=2$. If $p\neq 2$, we observe that for all $x\neq 0$,
$\phi(x)$ $:=\nabla\left(\frac{\|x\|_{p}^{2}}{2}\right)=\frac{1}{2}\nabla\Big{(}(\sum_{i}|x_{i}|^{p})^{2/p}\Big{)}=\frac{1}{2}\cdot\frac{2}{p\|x\|_{p}^{p-2}}\begin{bmatrix}\vdots\\ px_{i}|x_{i}|^{p-2}\\ \vdots\end{bmatrix}=\frac{1}{\|x\|_{p}^{p-2}}\begin{bmatrix}\vdots\\ x_{i}|x_{i}|^{p-2}\\ \vdots\end{bmatrix}.$
As
$\langle\phi(x),x\rangle=\frac{\|x\|_{p}^{p}}{\|x\|_{p}^{p-2}}=\|x\|_{p}^{2}$
and
$\|\phi(x)\|_{\ast}=\|\phi(x)\|_{p_{\ast}}=\frac{\Big{(}\sum_{i=1}^{d}|x_{i}|^{p_{\ast}(p-1)}\Big{)}^{1/p_{\ast}}}{\|x\|_{p}^{p-2}}=\frac{\|x\|_{p_{\ast}(p-1)}^{\frac{p_{\ast}(p-1)}{p_{\ast}}}}{\|x\|_{p}^{p-2}}=\frac{\|x\|_{p_{\ast}(p-1)}^{p-1}}{\|x\|_{p}^{p-2}}=\|x\|_{p},$
where the last equality follows as $p=p_{\ast}(p-1)$, $\|\cdot\|_{p}$ fulfills Definition 2.7 with $C=1$. ∎
Next, we prove regularity for the matrix norms known as Schatten $p$-norms. Recall that, for $p\in[1,\infty]$, the Schatten $p$-norm is defined for any matrix $A\in\mathbb{R}^{n\times m}$ with singular value decomposition $A=\sum_{i}\sigma_{i}u_{i}v_{i}^{\intercal}$ by the $p$-norm of the singular values, i.e.,
$\left\lVert A\right\rVert_{p}~{}:=~{}\Big{(}\sum_{i}\sigma_{i}^{p}\Big{)}^{1/p}.$
The special cases $p=1$, $p=2$ and $p=\infty$ are called the nuclear norm, the Frobenius norm and the spectral norm, respectively.
Lemma A.2.
For $p>1$, the Schatten $p$-norm is $1$-regular.
Proof.
By using, that
$\left\lVert A\right\rVert_{p}~{}:=~{}\Big{(}\sum_{i}\sigma_{i}^{p}\Big{)}^{1/p}=\Big{|}\mathrm{Tr}\left(\left(\sqrt{AA^{T}}\right)^{p}\right)\Big{|}^{1/p}$
by definition, together with the rules of differentiating matrices from Chapter 2.4 in [29], we find
$\phi(A)~{}=~{}\nabla\left(\frac{\left\lVert A\right\rVert_{p}^{2}}{2}\right)~{}=~{}\left\lVert A\right\rVert_{p}^{2-p}\sum_{i}\sigma_{i}^{p-1}u_{i}v_{i}^{\intercal}.$
Hence the inner product satisfies
$\langle\phi(A),A\rangle~{}=~{}\operatorname{tr}\left(\left\lVert A\right\rVert_{p}^{2-p}\sum_{i}\sigma_{i}^{p-1}v_{i}u_{i}^{\intercal}A\right)~{}=~{}\left\lVert A\right\rVert_{p}^{2-p}\operatorname{tr}\left(\sum_{i}\sigma_{i}^{p}v_{i}v_{i}^{\intercal}\right)~{}=~{}\left\lVert A\right\rVert_{p}^{2},$
while also, with $11/p+1/p_{*}=$ 1, the dual norm satisfies
$\left\lVert\phi_{p}(A)\right\rVert_{p_{*}}~{}=~{}\left\lVert A\right\rVert_{p}^{2-p}\left\lVert\sum_{i}\sigma_{i}^{p-1}u_{i}v_{i}^{\intercal}\right\rVert_{p_{*}}~{}=~{}\left\lVert A\right\rVert_{p}^{2-p+p/p_{*}}~{}=~{}\left\lVert A\right\rVert_{p}$
as required. ∎
Lemma A.3.
Let norm $\|\cdot\|_{a}$ defined on $\mathbb{R}^{d_{a}}$ be $C_{a}$-regular and norm $\|\cdot\|_{b}$ defined on $\mathbb{R}^{d_{b}}$ be $C_{b}$-regular. For $\lambda\in(0,1)$, we define the composite norm $\|\cdot\|_{a\circ b}$ as
$\|x\|_{a\circ b}=\left\|\begin{bmatrix}x_{a}\\ x_{b}\end{bmatrix}\right\|_{a\circ b}:=\sqrt{\lambda\|x_{a}\|_{a}^{2}+(1-\lambda)\|x_{b}\|_{b}^{2}},$
where $x_{a}:=x_{1:d_{a}}$ and $x_{b}:=x_{(d_{a}+1):d}$. Then the composite norm $\|\cdot\|_{a\circ b}$ is $\max\{C_{a},C_{b}\}$-regular.
Proof.
Set $\phi_{k}(y):=\nabla\left(\frac{\|y\|_{k}^{2}}{2}\right)$ and abbreviate $y=\begin{bmatrix}y_{1:d_{a}}\\ y_{(d_{a}+1):d}\end{bmatrix}:=\begin{bmatrix}y_{a}\\ y_{b}\end{bmatrix}$. Then
$\phi_{a\circ b}(x)$ $=\nabla\left(\frac{\|x\|_{a\circ b}}{2}\right)=\frac{1}{2}\nabla\left(\lambda\|x_{a}\|_{a}^{2}+(1-\lambda)\|x_{b}\|_{b}^{2}\right)$
$=\frac{1}{2}\begin{bmatrix}\lambda\nabla(\|x_{a}\|_{a}^{2})\\ (1-\lambda)\nabla(\|x_{b}\|_{b}^{2})\end{bmatrix}=\begin{bmatrix}\lambda\phi_{a}(x_{a})\\ (1-\lambda)\phi_{b}(x_{b})\end{bmatrix}.$
As $\|\cdot\|_{a}$ is $C_{a}$-regular and $\|\cdot\|_{b}$ $C_{b}$-regular,
$\langle\phi_{a\circ b}(x),x\rangle$ $=\lambda\langle\phi(x_{1:d_{a}}),x_{1:d_{a}}\rangle+(1-\lambda)\langle\phi(x_{(d_{a}+1):d}),x_{(d_{a}+1):d}\rangle$
$\geq\lambda\|x_{1:d_{a}}\|_{a}^{2}+(1-\lambda)\|x_{(d_{a}+1):d}\|_{b}^{2}=\|x\|_{a\circ b}^{2}.$
Further it follows, that
$\|\phi(x)\|_{\ast}$ $=\sup_{\|z\|\leq 1}\langle z,\phi(x)\rangle=\sup_{\lambda\|z_{a}\|_{a}^{2}+(1-\lambda)\|z_{b}\|_{b}^{2}\leq 1}\langle\begin{bmatrix}z_{a}\\ z_{b}\end{bmatrix},\begin{bmatrix}\lambda\phi_{a}(x_{a})\\ (1-\lambda)\phi_{b}(x_{b})\end{bmatrix}\rangle$
$=\sup_{\alpha\in[0,1]}\left(\sup_{\lambda\|z_{a}\|_{a}^{2}\leq\alpha}\lambda\langle z_{a},\phi_{a}(x_{a})\rangle+\sup_{(1-\lambda)\|z_{b}\|_{b}^{2}\leq 1-\alpha}(1-\lambda)\langle z_{b},\phi_{b}(x_{b})\rangle\right)$
$=\sup_{\alpha\in[0,1]}\left(\sqrt{\alpha\lambda}\sup_{\|z_{a}\|_{a}^{2}\leq 1}\langle z_{a},\phi_{a}(x_{a})\rangle+\sqrt{(1-\alpha)(1-\lambda)}\sup_{\|z_{b}\|_{b}^{2}\leq 1}\langle z_{b},\phi_{b}(x_{b})\rangle\right)$
$=\sup_{\alpha\in[0,1]}\left(\sqrt{\alpha\lambda}\|\phi_{a}(x_{a})\|_{a,\ast}+\sqrt{(1-\alpha)(1-\lambda)}\|\phi_{b}(x_{b})\|_{b,\ast}\right)$
$\leq\sup_{\alpha\in[0,1]}\left(\sqrt{\alpha\lambda}C_{a}\|x_{a}\|_{a}+\sqrt{(1-\alpha)(1-\lambda)}C_{b}\|x_{b}\|_{b}\right)$
$=\sqrt{\lambda C_{a}^{2}\|x_{a}\|_{a}^{2}+(1-\lambda)C_{b}^{2}\|x_{b}\|_{b}^{2}}$
$\leq\max\{C_{a},C_{b}\}\left\lVert x\right\rVert_{a\circ b}.$
This shows that $\|\cdot\|_{a\circ b}$ is $C$-regular with $C=\max\{C_{a},C_{b}\}$ according to Definition 2.7. ∎
Lemma A.4.
For $s\geq 1$, we define $\phi_{s}(x):=\nabla\Big{(}\frac{\|x\|^{s}}{s}\Big{)}.$
[…]
$\begin{cases}\langle\phi_{s}(x),x\rangle&\geq\|x\|^{s}\\ \|\phi_{s}(x)\|_{\ast}^{\alpha}&\leq C^{\alpha}\|x\|^{\alpha(s-1)}.\end{cases}$
@@
$\langle\phi_{s}(x),x\rangle$ $=\frac{\langle\phi(x),x\rangle}{\|x\|^{2-s}}\geq\frac{\|x\|^{2}}{\|x\|^{2-s}}=\|x\|^{s}$
$\|\phi_{s}(x)\|_{\ast}^{\alpha}$ $=\frac{\|\phi(x)\|_{\ast}^{\alpha}}{\|x\|^{\alpha(2-s)}}\leq\frac{C^{\alpha}\|x\|^{\alpha}}{\|x\|^{\alpha(2-s)}}=C^{\alpha}\|x\|^{\alpha(s-1)},$

### v2

@@ 4 Binary search for Hölder smooth functions
In each round $t$, Algorithm 1 performs binary search to find the momentum parameter $\lambda_{t}$ that satisfies
$\lambda_{t}g^{\prime}(\lambda_{t})+C_{t}g(\lambda_{t})\leq\epsilon_{t},$ (6)
where we use the abbreviation $g(\lambda)\triangleq F(\lambda x_{t}^{ag}+(1-\lambda)x_{t})-F(x_{t}^{ag})$ so that in particular $g(1)=0$. The condition (6) is crucial for the convergence proof of Theorem 3.1. The core idea behind our Algorithm 2 is inspired by Algorithm 2 of [pmlr-v125-hinder20a]. We aim to efficiently find a $\lambda_{t}$ that satisfies (6) in the more general setting where $F$ is weakly smooth with respect to a possibly non-Euclidean norm.
Note that if $g(0)\leq 0$ or $g$ is decreasing at $\lambda=1$, then (6) is immediately satisfied. If neither of these conditions hold, then $g(0)>g(1)$ and $g^{\prime}(1)>0$, indicating that $g$ must switch at least once from decreasing to increasing. For the largest $\lambda_{*}$ fulfilling $g^{\prime}(\lambda_{*})=0$ it holds that $g(\lambda_{*})<0$. We further show in the proof of Theorem 4.1 that all $\lambda\in[\lambda_{*}-\delta,\lambda_{*}]$, with a properly chosen $\delta$, must satisfy condition (6). The efficiency of our algorithm in finding any $\lambda$ in that interval is stated in Theorem 4.1 below.
Algorithm 2 Generalized binary search
@@
  if $g^{\prime}(\lambda_{t})+C_{t}g(\lambda_{t})\leq\epsilon_{t}$
@@
Under Assumption 1 (in particular $F$ is $(L,\kappa)$-weakly smooth), and for fixed stopping parameters $(C_{t},\epsilon_{t})$ and endpoints $(x_{t},x_{t}^{ag})$, Algorithm 2 finds $\lambda$, satisfying
@@
[pmlr-v125-hinder20a] analyze their binary search algorithm for smooth functions, i.e., specifically with $\kappa=2$ in Definition 2.3. Our algorithm accommodates a general $\kappa$, addressing the broader class of weakly smooth functions. While the running time depends on $\kappa$, it appears only as an exponent in the logarithm. Consequently, this introduces only a constant multiplicative factor, thereby generalizing the algorithm of [pmlr-v125-hinder20a].
@@
Proof Sketch. The proof can be roughly divided into three parts. We first show that there exists a $\lambda_{*}$ satisfying the strengthening of condition (6) with $\epsilon_{t}=0$. Next, we show that this implies that condition (6) holds for all $\lambda\in[\lambda_{*}-\delta,\lambda_{*}]$, where $\delta$ is properly chosen depending on the weak smoothness parameters $(L,\kappa)$ of $F$ and the parameters $(C_{t},\epsilon_{t},x_{t},x_{t}^{ag})$ of the algorithm. Finally we prove that our algorithm finds a $\lambda$ within that interval after at most $\mathcal{O}\Big{(}\log\Big{(}\frac{LC_{t}\|x_{t}-x_{t}^{ag}\|^{\kappa}}{\epsilon_{t}}\Big{)}\Big{)}$ iterations (details in Appendix E).
4.1 Polynomial growth of our iterates
To complete our analysis, we present upper bounds for the iterates $(x_{t},x_{t}^{ag})$ generated by Algorithm 1. In fact, our Theorem 4.1 shows that the number of iterations is upper bounded by a term logarithmic in $t$ provided the distance $\|x_{t}-x_{t}^{ag}\|$ grows at most polynomially in $t$. This result is formalized in the following theorem.
Theorem 4.2.
Let $D_{\psi}(x_{\star},x_{1})\leq B$. Under Assumption 1, suppose we are running Algorithm 1 such that for all $t\geq 1$
$\max\left(\Big{(}\frac{qL\eta_{t}}{\kappa}\Big{)}^{\frac{1}{q-1}},\Big{(}\frac{L\alpha_{t}}{\kappa}\Big{)}^{\frac{1}{q-1}}\right)\leq K^{n_{1}}t^{n_{2}},$
where $n_{1},n_{2}\geq 0$, and $K^{n_{1}}>B$ with $K^{n_{1}}$ potentially depending on $(L,\tau,\mu)$. Define the iterate radius by $R_{t}=\max\{\|x_{t}-x_{\star}\|,\|x_{t}^{ag}-x_{\star}\|,\|x_{t}^{md}-x_{\star}\|\}$. Then, for all $t\geq 1$, we have
$R_{t}=\mathcal{O}\Big{(}K^{\frac{n_{1}(q-1)}{q-\kappa}}t^{\frac{(q-1)(n_{2}+1)}{q-\kappa}}\Big{)}.$
4.2 Lower bound for Binary search
Although the following theorem does not contribute to our upper bounds, we analyze the binary search algorithm without assuming star-convexity to highlight the efficiency of Algorithm 2 under weak smoothness. This section shows that the number of iterations needed in Theorem 4.1 to find a suitable $\lambda$ satisfying condition (6) is both sufficient and necessary - fewer itations could yield in a $\lambda$ that does not satisfy condition (6). To show this lower bound, we determine the minimum number of iteration steps required to find a suitable $\lambda$ for a univariate smooth target function.
Theorem 4.3.
Let $C,\epsilon,L_{\star}>0$. Consider a sequence of $N$ points $(\lambda_{1},\ldots,\lambda_{N})$ evaluated at each iteration $(g(\lambda_{i}),g^{\prime}(\lambda_{i}))$. If the number of iterations is insufficient meaning
$N<\log(5)\Big{(}\log\frac{L_{\star}}{\epsilon}+\log\frac{C}{(C+1)^{2}}-\log(88)\Big{)},$
no algorithm can find $\lambda$ fulfilling condition (6).
The theorem shows that if $N\lesssim\log(L_{\star}/\epsilon)$, it becomes impossible to determine $\lambda$ with precision $\epsilon$. The full version of this lower bound is given in Theorem F.5.
5 Summary
By combining the convergence results from Corollary 3.2 with the complexity analysis presented in Theorem 4.1, we establish a bound on the precision error of Algorithm 1 in terms of the number of first-order oracle calls.
Corollary 5.1.
Let $T>0$ be the number of iterations and $\frac{1}{\mu}D_{\psi}(x_{\star},x_{1})\leq B$. The tuning of Theorem 3.1 with $\alpha~{}=~{}\frac{\mu}{L}\left(\frac{(q-\kappa)B}{\kappa}\right)^{\frac{q-\kappa}{q}}$ guarantees $F(x_{T+1}^{ag})-F_{\star}~{}=~{}\mathcal{O}_{q,\kappa}\left(L\tau^{q}B^{\frac{\kappa}{q}}\frac{\log(T)}{T^{\frac{\kappa q+\kappa-q}{q}}}\right).$ The algorithm’s oracle usage over $T$ iterations is upper bounded by $\mathcal{O}(T\log(LB\tau T))$, where this bound represents the maximum number of times the oracle is called during the entire execution process.
To apply Corollary 5.1 to the $p$-norms in Example 3.1, we consider a problem where both the initial points $x_{1}$ and minimizer $x_{\star}$ lie within a $p$-norm ball of radius $R$, i.e.,  $\left\lVert x_{1}\right\rVert_{p},\left\lVert x_{\star}\right\rVert_{p}\leq R$, which implies that $\left\lVert x_{1}-x_{\star}\right\rVert_{p}\leq 2R$ by the triangle inequality. For $p$-norms, recall that we use the distance-generating function $\psi(\cdot)=\frac{1}{q}\left\lVert\cdot\right\rVert_{p}^{q}$ with $q=\max\{2,p\}$. In Appendix H, we show that the Bregman divergence also satisfies $D_{\psi}(x_{\star},x_{1})\leq 2R^{q}$. Combining Theorems 3.1 and 4.2, we obtain the following convergence guarantees.
Corollary 5.2.
Consider the setting of Assumption 1 and let $\left\lVert\cdot\right\rVert_{p}$ denote the $p$-norm.
  • (i)
    For $1<p\leq 2$ and $\kappa<2$, Algorithm 1 with $q=2$ returns an $\epsilon$-optimal solution within
    $\mathcal{O}_{\kappa}\left(\left(\frac{L\tau^{2}R^{\kappa}}{\epsilon}\right)^{\frac{2}{3\kappa-2}}\log^{2}\Big{(}\frac{L\tau R}{\epsilon}\Big{)}\right)$
    evaluations of the function value and gradient.
  • (ii)
    For $p>2$, Algorithm 1 with $q=p$ returns an $\epsilon$-optimal solution within
    $\mathcal{O}_{p,\kappa}\left(\Big{(}\frac{L\tau^{p}R^{\kappa}}{\epsilon}\Big{)}^{\frac{p}{\kappa p+\kappa-p}}\log^{2}\Big{(}\frac{L\tau R}{\epsilon}\Big{)}\right)$
    evaluations of the function values and the gradients.
Note that the convergence guarantees of Corollary 5.2 for star-convex functions match the lower bounds proved for convex functions, completing Table 1. For $p=1$, our analysis applies with a dimension factor, detailed in Appendix I. For the smooth case $\kappa=2$ with $21<p<$ 2, where iterates may grow exponentially, we extend our algorithm while maintaining similar complexity; further details are provided in Appendix J.
@@ 6 Conclusion and future research
In this work, we introduce a novel class of structured non-convex functions, namely $\tau$-star-convex functions that are weakly smooth with respect to an arbitrarily norm. Our framework applies to a broad subclass of star-convex functions and generalizes the setting analyzed in [pmlr-v125-hinder20a] which focuses on smooth star-convex functions with respect to the Euclidean norm.
We develop the accelerated Algorithm 1 to efficiently minimize any weakly smooth function within this broad class, building on and extending the binary search technique from [pmlr-v125-hinder20a]. Our analysis shows that the algorithm achieves a near-optimal accelerated convergence rate for all $p$-norms with $p>1$. Additionally, our algorithm has a runtime dependence on $\tau^{\max\{2,p\}}$ for $p\geq 1$, shown to be optimal in the Euclidean case when $p=2$, see [pmlr-v125-hinder20a].
While our work focuses on unconstrained minimization problems, future research could extend our approach to constrained settings. Currently, the proximal step for calculating $x_{t+1}^{ag}$ in Algorithm 1 is shown to be effective only for unconstrained problems. For the constrained non-convex case, since evaluating the direction of the constrained proximal step becomes a challenge, it is unknown if accelerated convergence rates are possible.
Finally, our method operates within a first-order access model, meaning that the algorithm has access to both the objective function value and gradient at the evaluation points. It remains an open problem whether this can be extended to “pure” first-order access models, that rely solely on gradient evaluations. While this extension is feasible in the convex case, it is unclear whether it can be achieved for star-convex functions. Another promising direction is to explore the applicability of our method in a stochastic model. Although stochastic accelerated mirror descent has been previously analyzed in the literature [NIPS2009_ec5aa0b7], the main challenge lies in adapting the binary search algorithm to the noisy setting.
[…]
Appendix A A regularity property of norms
[…]
$\phi(x)~{}:=~{}\nabla\left(\frac{\|x\|^{2}}{2}\right)~{}=~{}\operatorname*{argmax}_{y:\|y\|_{\ast}=\|x\|}~{}y^{\intercal}x.$
Wherever the norm is not differentiable, we may pick $\phi(x)$ to be any sub-gradient or maximizer. To see the equality, recall that $\|x\|=\max_{z:\|z\|_{\ast}=1}z^{\intercal}x$ and let $z$ attain that maximum. Then differentiation [^1] Recall that a sub-gradient of a maximum $x\mapsto\max_{y}f(x,y)$ is the derivative of the objective $\nabla_{x}f(x,y_{*})$ with a maximizer $y^{*}$ held fixed. gives $\phi(x)=\|x\|z$. Reparametrising by $y=\|x\|z$ gives the right hand side.
Fact A.1 (Regularity of a norm).
For all $x\in\mathbb{R}^{d}$
$\langle\phi(x),x\rangle$ $~{}=~{}\|x\|^{2},$
$\|\phi(x)\|_{\ast}$ $~{}=~{}\|x\|.$
These two properties are crucial in our accelerated framework (Algorithm 1), especially for weakly smooth objective functions. In particular, they allow for the choice of aggressive step-sizes and the derivation of accelerated convergence rates.
@@ Appendix A A regularity property of norms
For $s\geq 1$, we define $\phi_{s}(x):=\nabla\Big{(}\frac{\|x\|^{s}}{s}\Big{)}.$ For all $\alpha\geq 0$
$\begin{cases}\langle\phi_{s}(x),x\rangle&=\|x\|^{s}\\ \|\phi_{s}(x)\|_{\ast}^{\alpha}&=\|x\|^{\alpha(s-1)}.\end{cases}$
@@
$\langle\phi_{s}(x),x\rangle$ $=\frac{\langle\phi(x),x\rangle}{\|x\|^{2-s}}=\frac{\|x\|^{2}}{\|x\|^{2-s}}=\|x\|^{s},$
$\|\phi_{s}(x)\|_{\ast}^{\alpha}$ $=\frac{\|\phi(x)\|_{\ast}^{\alpha}}{\|x\|^{\alpha(2-s)}}=\frac{\|x\|^{\alpha}}{\|x\|^{\alpha(2-s)}}=\|x\|^{\alpha(s-1)},$